# Jet-like structures in low-mass binary neutron star merger remnants

### v1

@@ II Methods
To perform the numerical simulations we use the in-house and well established Illinois GRMHD code Etienne et al. (2010, 2012a); Farris et al. (2012) and the methods described in our previous works (see e.g. Ruiz et al. (2021, 2020a, 2019); Ruiz and Shapiro (2017)). Illinois GRMHD uses the Carpet code Goodale et al. (2003); Schnetter et al. (2004) for moving-box mesh refinement. We use the Baumgarte–Shapiro–Shibata–Nakamura (BSSN) formulation of the Einstein equations Shibata and Nakamura (1995); Baumgarte and Shapiro (1998) with the moving-puncture gauge condition (Eqs. (2)-(4) in Etienne et al. (2008)) with the damping parameter $\eta$ in the shift evolution equation set to $\eta\approx 2.0/M$, where $M$ is the total ADM mass of the system. At the boundaries we apply outgoing-wave or Sommerfeld boundary conditions to all the BSSN variables. We use fourth order centered stencils for spatial derivatives, except for shift advection terms where fourth order upwind stencils are used. Time integration is performed using the Method of Lines with a fourth order Runge-Kutta integration scheme with a Courant- Friedrichs-Lewy (CFL) factor set to 0.45. To control spurious high frequency noise fifth order Kreiss-Oliger dissipation Baker et al. (2006) is added to the evolution equations.
For the matter evolution we evolve the equations of ideal MHD in conservative form using a high-resolution shock capturing method (see Eqs. (27)-(29) in Etienne et al. (2010)) which employs the piecewise parabolic reconstruction scheme (PPM) Colella and Woodward (1984) and the Harten, Lax, and van Leer (HLL) approximate Riemann solver Harten et al. (1983). We evolve the magnetic field by integrating the magnetic induction equation using a vector potential in order to ensure it remains divergenceless throughout the evolution (see Eqs. (19)-(20) in Etienne et al. (2010)). We also use the generalized Lorentz gauge in Farris et al. (2012) to avoid the build-up of superious magnetic fields Etienne et al. (2012a) with a damping factor of $\chi=2.0/M$.
@@ II.1 Initial data
$A_{\phi}=\frac{\pi\varpi^{2}I_{0}r^{2}_{0}}{(r^{2}_{0}+r^{2})^{3/2}}\left[1+\frac{15r^{2}_{0}(r^{2}_{0}+\varpi^{2})}{8(r^{2}_{0}+r^{2})^{2}}\right],$ (2)
@@
which approximately corresponds to that generated by a current loop inside the NS with radius $r_{0}$ and current $I_{0}$ (Fig. 2 upper left). Here $r^{2}=\varpi^{2}+z^{2}$ and $\varpi^{2}=(x-x_{\rm NS})^{2}+(y-y_{\rm NS})^{2}$, where the center of the star $(x_{\rm NS},y_{\rm NS},0)$ is determined by the coordinate of maximum rest-mass density. We set $I_{0}$ and $r_{0}$ such that the maximum value of the magnetic-to-gas-pressure ratio in the NS interior is $\beta^{-1}:=P_{B}/P_{\textup{gas}}=0.0023$, which corresponds to a maximum magnetic field as measured by a normal observer of $B_{\textup{max}}\sim 10^{15.2}$ at the pole of the NS. This is significantly larger than the surface magnetic field strengths of $10^{8}-10^{12.2}$ expected for NSs in binary systems as inferred from observations of binary pulsars Tauris et al. (2017); Lorimer (2008). However, this value is chosen to model the field strengths expected due to the exponential growth from magnetic instabilities, initially driven by the Kelvin-Helmholtz Instability (KHI) at the shearing surface when the stars first collide. This instability boosts the magnetic energy by almost a factor of 10 within several ms, until the instability saturates or the shear surface is destroyed by shocks Kiuchi et al. (2015). The magnetic energy growth rate in the linear regime is inversely proportional to the minimum resolvable wavelength, and thus to the numerical resolution, at least down to a resolution $\Delta x_{\textup{min}}\sim 12.5$m Price and Rosswog (2006); Kiuchi et al. (2024). The magnetorotational instability (MRI), magnetic winding and potentially the Rayleigh-Taylor instability also work to boost the magnetic field. Detailed special relativistic and approximate-GR simulations have shown that the field can be amplified to magnetar levels of $\gtrsim 10^{15}$ as the stretching and folding of the magnetic field lines converts kinetic energy to magnetic Price and Rosswog (2006); Rosswog (2007); Obergaulinger et al. (2010); Oechslin et al. (2007); Zrake and MacFadyen (2013). Very high resolution GRMHD simulations have shown similar results, with the $\Delta x_{\textup{min}}=37$m simulation of Aguilera-Miret et al. Aguilera-Miret et al. (2020) showing amplification from $10^{11}$ to $10^{16}$ within the first 5ms post-merger, and the $\Delta x_{\textup{min}}=17.5$m simulation of Kiuchi et al. (2015) Kiuchi et al. (2015) where an initial field strength of $10^{13}$ is amplified to create a large-scale $\gtrsim 10^{16}$ poloidal field due to a combination of the KHI and the MRI. Computational cost limits the resolution we can use here, so the KHI growth rate in our work is consequently suppressed compared to expected physical reality. Following our previous works Ruiz et al. (2016); Ruiz and Shapiro (2017); Ruiz et al. (2021) we therefore adopt the artificially strong initial magnetic field to compensate, as is common practice in GRMHD simulations of NSNS mergers Ciolfi et al. (2019). Notice that this magnetic field modifies the pressure of the system in $<1\%$ and hence it does not have a significant impact on the NS structure or the inspiral phase of the binary. During this phase the magnetic field is simply advected by the fluid. An alternative approach using large-eddy simulations and subgrid-scale models is discussed in Giacomazzo et al. (2015); Aguilera-Miret et al. (2020); Palenzuela et al. (2022); Aguilera-Miret et al. (2022, 2023); Izquierdo et al. (2024). The idea is to compensate for the limited resolution by including additional terms into the evolution equations that attempt to capture the otherwise unresolved subgrid dynamics. The advantage is that it allows you to start with smaller magnetic fields, more similar to those of known pulsars, and it is argued that it is better able to model the small scale quasi-isotropic turbulent field generated via the KHI Aguilera-Miret et al. (2020). The downside is that the results may depend heavily on the choice of subgrid model and the coefficients chosen for the subgrid terms.
@@ II.3 Diagnostics
Figure 3: Conservation of the mass and angular momentum integrals $M_{\textup{int}}$ and $J_{\textup{int}}$ for the representative SP2.54 case vs, retarded time $t_{\textup{ret}}:=t-r_{\textup{ext}}$. The magenta lines denote the change in the surface integrals Eqs. (37) and (39) in Etienne et al. (2012a) evaluated on spherical surfaces at $r_{\textup{ext}}\simeq 273M$. The cyan line denotes the total EM energy (upper plot) and $z$ component of the total EM angular momentum (lower plot) that has passed across that surface. The green and red lines do the same for the fluid and GWs respectively. The black dashed line gives the total energy in the top panel, and total angular momentum in the bottom panel that should be conserved, adding in the lost energy and angular momentum.
@@ II.4 Criteria for a sGRB-compatible jet
As discussed in Sec.I, to be consistent with sGRB observations, the central engine needs to produce a an ultrarelativistic collimated jet with a maximum Lorentz factor of $\Gamma\gtrsim 20$ and, to match the kilonova observations, an ejecta mass of $\gtrsim 1\%$ of the total NSNS mass Coulter et al. (2017); Shappee et al. (2017); Arcavi et al. (2017). One highly non-trivial challenge in analyzing the result of GRMHD merger simulations is estimating the terminal Lorentz factor $\Gamma_{\infty}$, as our simulations have a finite spatial and temporal extent while we expect the acceleration to increase with distance from the central engine (for instance, for a simple model of an ideal force-free paraboloidal jet the maximum Lorentz factor goes with distance along the jet $z$ as $\Gamma_{\textup{max}}\sim 0.3(z/M)^{1/2}$ Tchekhovskoy et al. (2008); McKinney and Blandford (2009); Narayan et al. (2010)). The physics of such jets, whether in the context of sGRB progenitors or AGN, has been explored extensively through both numerical MHD and GRMHD studies (see e.g. Tchekhovskoy et al. (2008); McKinney and Blandford (2009); Narayan et al. (2010); Fernández et al. (2019); Nathanail et al. (2021); Gottlieb et al. (2022a); Mizuno (2022); Perucho and López-Miralles (2023) and the references cited therein) and semi-analytic studies (see e.g. Vlahakis and Konigl (2003); Beskin and Nokhrina (2005); Komissarov (2011); Lyubarsky (2011)) which make use of the assumptions of ideal MHD, special relativity, axisymmetry, steady state and radial self-similarity. For a steady state, axisymmetric ideal MHD flow integrating the equations of motion results in a number of conserved quantities along the poloidal component $\bm{B}_{p}$ of the magnetic field lines, which are parallel to the poloidal component of the fluid velocity $\bm{v}_{p}$ Vlahakis and Konigl (2003); Narayan et al. (2010); Lyubarsky (2011). One of these quantities is the ratio of total energy flux to rest-mass flux along a bundle of field lines, given by
$\mu:=$ energy fluxrest-mass $\frac{\textup{energy flux}}{\textup{rest-mass flux}}=\;\frac{\Gamma^{2}\rho_{0}hv_{p}+\tfrac{1}{4\pi}(\bm{E}\times\bm{B})\cdot\hat{\bm{v}}_{p}}{\Gamma\rho_{0}v_{p}}$
@@
Figure 4: Top panel: Pseudocolor plot of the quality factor $Q\equiv\lambda_{\textup{MRI}}/dx$ on the equatorial plane. Bottom panel: Rest-mass density normalised to its initial maximum value on a log scale along with $\lambda_{\textup{MRI}}$ (white line) on the meridional plane. The height of the line above the $x$ axis indicates the value of $\lambda_{\textup{MRI}}$ at $x=0$. Note that almost all of the area in the top image is red, indicating that the quality factor is $Q\gtrsim 10$ as required. We see that the height of the remnant and the height of the surrounding torus is larger than $\lambda_{\textup{MRI}}$ wavelength in all but a few regions, suggesting the MRI is active. The data shown here is for the SP2.57 case depicted at $\sim 17.8$ms post merger. However, it is representative for all cases, including the HMNS remnant in case IR2.70 prior to its collapse.
@@ II.5.4 $\alpha\Omega$ dynamo
Figure 5: 2D plots in the meridional plane at 20ms post-merger for five of the simulations in Table 2. We show the rest-mass density $\rho_{0}$ on a log scale relative to the initial maximum rest-mass density (top row), $b^{2}/(2\rho_{0})$ which approximately corresponds to the magnetization $\sigma$ (second row), the Lorentz factor (third row), and the strength of the poloidal and toroidal magnetic field (fourth and fifth rows), with the magnetic field lines shown in the poloidal field plots. The scale on the two axes is the same. The arrows in the Lorentz factor plot (third row) indicate the flow velocities, and the arrow labelled on the left hand side indicates the magnitude of the speed of light.
@@ III Results
Figure 7: The accretion rate onto the BH remnant for the IR2.70 case (top), the fraction of the total rest mass outside the apparent horizon (middle), and the jet launching efficiency ($\eta=L/\dot{M}$ where $L=L_{\textup{EM}}+L_{\textup{fluid}}$ is the jet luminosity) on log scales. The accretion rate decays as $\dot{M}\propto t^{-2}$ in the first $\sim 0.5$ms post-BH formation, as expected for highly magnetized disks Gottlieb et al. (2023), then rapidly transitions to $\dot{M}\propto t^{-1}$ as for hydrodynamically-dominated accretion, where accretion is driven by shocks and spirals in the disk Gottlieb et al. (2023). We also show the expected accretion efficiency for a Novikov-Thorne thin disk model (green dashed line, lower panel).
@@
Accretion into the BH down the polar axis and the acceleration of gas outwards due to the magnetic forces produces a low density ($\rho_{0}\;\lesssim 10^{7}\textup{g}\,\textup{cm}^{-3}$) evacuated funnel with a half-opening angle of $\sim 30^{\circ}$ (see the top left image of Fig. 5). [^5] Notice that in our previous NSNS merger simulations where neutrinos were incorporated using the truncated-moment (M1) formalism the baryon pollution in the funnel was a factor of $\sim 10$ lighter than in this case Sun et al. (2022). The low density and strong $\sim 10^{16}$ magnetic field enables a force-free environment with $b^{2}/(2\rho_{0})\sim\sigma\gg 1$ in the funnel above the BH poles, and the helical field enables the acceleration of an unbound, collimated, mildly relativistic ($\Gamma\sim 2.0$ in our simulation box) outflow, mainly along the outer regions of the funnel. This behaviour is consistent with the standard picture of Poynting-flux driven outflow reported in the literature (e.g. McKinney and Blandford (2009)) which we describe as an incipient jet. The accretion rate (see Fig. 7 top panel) shows a power-law decay with time, initially falling with $\dot{M}\propto t^{-2}$ in the first $\sim 0.5$ms, consistent with the expectation for a magnetically dominated disk Gottlieb et al. (2023), then rapidly transitions to $\dot{M}\propto t^{-1}$ for the remaining simulation time, consistent with a hydrodynamically-dominated disk where accretion is driven by shocks and density spirals Gottlieb et al. (2023). The accretion efficiency (see Fig. 7 bottom panel), $\eta:=(L_{\textup{fluid}}+L_{\textup{EM}})/\dot{M}$, increases to $\sim 0.09$, close to the value expected for the thin disk Novikov-Thorne model (see Shapiro (2005)), which has shown remarkably close agreement with numerical MHD simulations of BZ jets McKinney and Gammie (2004); Shapiro (2005), but below the maximum value for a magnetically-arrested disk (MAD) Lowell et al. (2024). From the $\dot{M}\propto t^{-1}$ accretion rate we can estimate the lifetime of the accretion disk as $\sim 2.2$s, very close to the typical lifetime of a sGRB, although our lack of knowledge of the precise details of how the energy from the incipient jet is converted to the $\gamma$-ray signal we observe limits the usefulness of this comparison.
@@
where $B_{\textup{pol}}$ is the poloidal magnetic field measured at the BH pole. The luminosity from Eq. (23) is plotted as connected dots in the middle panel of Fig. 8. We can see that the EM luminosity for the IR2.70 case converges towards the $L_{\textup{BZ}}$ value towards the end of our simulation as the additional power from the magnetized torus dies down and the BZ mechanism dominates. The magnitude lies within the narrow “universal range” of luminosities shown in Shapiro (2017) to arise from BH + disk systems that result from compact binary mergers containing NSs, as well as from the magnetorotational collapse of massive stars. This, combined with the other criteria already validated for the very similar case reported in Ruiz et al. 2021 Ruiz et al. (2021), demonstrates that the BZ mechanism is operating in our simulation, as we concluded in Ruiz et al. (2016, 2018a); Paschalidis et al. (2015); Sun et al. (2022). Although we only observe a mildly relativistic outflow within our simulation region, an extensive literature suggests that the BZ effect alone can ultimately accelerate gas to ultrarelativistic ($\Gamma_{\infty}\gg 100$) velocities, even with significant baryon pollution (see e.g. Paschalidis et al. (2015); McKinney (2005)) and therefore provide the central engine for a sGRB-compatible jet.
Figure 8: Luminosity via fluid outflow, EM radiation and gravitational waves, vs retarted time since $t_{\textup{merge}}$ (defined as the time of peak GW emission) extracted over a sphere at $r_{\textup{ext}}=1024\textup{km}\sim 260M$. The purple dots and connecting line in the middle plot show the expected BZ electromagnetic luminosity from Eq. (23) for the BH mass, spin and approximate $B_{\textup{pol}}$ for case IR2.70.

### v2

@@ II Methods
To perform the numerical simulations we use the in-house and well established Illinois GRMHD code Etienne et al. (2010, 2012a); Farris et al. (2012) and the methods described in our previous works (see e.g. Ruiz et al. (2021, 2020a, 2019); Ruiz and Shapiro (2017)). Illinois GRMHD uses the Carpet code Goodale et al. (2003); Schnetter et al. (2004) for moving-box mesh refinement. We use the Baumgarte–Shapiro–Shibata–Nakamura (BSSN) formulation of the Einstein equations Shibata and Nakamura (1995); Baumgarte and Shapiro (1998) with the moving-puncture gauge condition (Eqs. (2)-(4) in Etienne et al. (2008)) with the damping parameter $\eta$ in the shift evolution equation set to $\eta\approx 2.0/M$, where $M$ is the total Arnowitt–Deser–Misner (ADM) mass of the system. At the boundaries we apply outgoing-wave or Sommerfeld boundary conditions to all the BSSN variables. We use fourth order centered stencils for spatial derivatives, except for shift advection terms where fourth order upwind stencils are used. Time integration is performed using the Method of Lines with a fourth order Runge-Kutta integration scheme with a Courant- Friedrichs-Lewy (CFL) factor set to 0.45. To control spurious high frequency noise fifth order Kreiss-Oliger dissipation Baker et al. (2006) is added to the evolution equations.
For the matter evolution we evolve the equations of ideal MHD in conservative form using a high-resolution shock capturing method (see Eqs. (27)-(29) in Etienne et al. (2010)) which employs the piecewise parabolic reconstruction scheme (PPM) Colella and Woodward (1984) and the Harten, Lax, and van Leer (HLL) approximate Riemann solver Harten et al. (1983). We evolve the magnetic field by integrating the magnetic induction equation using a vector potential in order to ensure it remains divergenceless throughout the evolution (see Eqs. (19)-(20) in Etienne et al. (2010)). We also use the generalized Lorentz gauge in Farris et al. (2012) to avoid the build-up of spurious magnetic fields Etienne et al. (2012a) with a damping factor of $\xi=2.0/M$.
@@ II.1 Initial data
$A_{\phi}=\frac{15\pi\varpi^{2}I_{0}r^{2}_{0}}{23(r^{2}_{0}+r^{2})^{3/2}}\left[1+\frac{15r^{2}_{0}(r^{2}_{0}+\varpi^{2})}{8(r^{2}_{0}+r^{2})^{2}}\right],$ (2)
@@
which approximately corresponds to that generated by a current loop inside the NS with radius $r_{0}$ and current $I_{0}$ (Fig. 2 upper left). Here $r^{2}=\varpi^{2}+z^{2}$ and $\varpi^{2}=(x-x_{\rm NS})^{2}+(y-y_{\rm NS})^{2}$, where the center of the star $(x_{\rm NS},y_{\rm NS},0)$ is determined by the coordinate of maximum rest-mass density. We set $I_{0}$ and $r_{0}$ such that magnetic field as measured by a normal observer at the NS pole is $B_{\textup{pole}}=10^{15.2}$ and the maximum value is $B_{\textup{max}}\sim 10^{16.7}$ in the NS center. The maximum value of the magnetic-to-gas-pressure ratio in the NS interior is $\beta^{-1}:=P_{B}/P_{\textup{gas}}=0.0023$. The magnetic field at the pole is significantly larger than the surface magnetic field strengths of $10^{8}-10^{12.2}$ expected for NSs in binary systems as inferred from observations of binary pulsars Tauris et al. (2017); Lorimer (2008). However, this value is chosen to model the field strengths expected due to the exponential growth from magnetic instabilities, initially driven by the Kelvin-Helmholtz Instability (KHI) at the shearing surface when the stars first collide. This instability boosts the magnetic energy by almost a factor of 10 within several ms, until the instability saturates or the shear surface is destroyed by shocks Kiuchi et al. (2015). The magnetic energy growth rate in the linear regime is inversely proportional to the minimum resolvable wavelength, and thus to the numerical resolution, at least down to a resolution $\Delta x_{\textup{min}}\sim 12.5$m Price and Rosswog (2006); Kiuchi et al. (2024). The magnetorotational instability (MRI), magnetic winding and potentially the Rayleigh-Taylor instability also work to boost the magnetic field. Detailed special relativistic and approximate-GR simulations have shown that the field can be amplified to magnetar levels of $\gtrsim 10^{15}$ as the stretching and folding of the magnetic field lines converts kinetic energy to magnetic Price and Rosswog (2006); Rosswog (2007); Obergaulinger et al. (2010); Oechslin et al. (2007); Zrake and MacFadyen (2013). Very high resolution GRMHD simulations have shown similar results, with the $\Delta x_{\textup{min}}=37$m simulation of Aguilera-Miret et al. Aguilera-Miret et al. (2020) showing amplification from $10^{11}$ to $10^{16}$ within the first 5ms post-merger, and the $\Delta x_{\textup{min}}=17.5$m simulation of Kiuchi et al. (2015) Kiuchi et al. (2015) where an initial field strength of $10^{13}$ is amplified to create a large-scale $\gtrsim 10^{16}$ poloidal field due to a combination of the KHI and the MRI. Computational cost limits the resolution we can use here, so the KHI growth rate in our work is consequently suppressed compared to expected physical reality. Following our previous works Ruiz et al. (2016); Ruiz and Shapiro (2017); Ruiz et al. (2021) we therefore adopt the artificially strong initial magnetic field to compensate, as is common practice in GRMHD simulations of NSNS mergers Ciolfi et al. (2019). Notice that this magnetic field modifies the pressure of the system in $<1\%$ and hence it does not have a significant impact on the NS structure or the inspiral phase of the binary. During this phase the magnetic field is simply advected by the fluid. While it has been suggested that purely poloidal fields are unstable on an Alfén time scale Ciolfi et al. (2011); Lasky et al. (2012), our previous works have shown that a strong, large-scale poloidal field is a requirement for jet launching (e.g. Etienne et al. (2012b); Ruiz et al. (2020a)), and this idealized topology both allows us to resolve the MRI instability (as the MRI wavelength is proportional to $|b^{P}|$) and serves as a useful model. Hence, while the true pre-merger magnetic field structure inside neutron stars remains uncertain Braithwaite and Spruit (2006); Bilous et al. (2019); Tsokaros et al. (2022) a strong poloidal field provides a “best-case” scenario for jet launching.
An alternative approach using large-eddy simulations and subgrid-scale models is discussed in Giacomazzo et al. (2015); Aguilera-Miret et al. (2020); Palenzuela et al. (2022); Aguilera-Miret et al. (2022, 2023); Izquierdo et al. (2024). The idea is to compensate for the limited resolution by including additional terms into the evolution equations that attempt to capture the otherwise unresolved subgrid dynamics. The advantage is that it allows you to start with smaller magnetic fields, closer to those of known pulsars, and it is suggested that it is better able to model the small scale quasi-isotropic turbulent field generated via the KHI Aguilera-Miret et al. (2020), and that it is preferable to using artificially strong initial poloidal fields which the authors in Aguilera-Miret et al. (2023) argue could lead to unrealistic outcomes. However, the downside is that the results may depend heavily on the choice of subgrid model and the coefficients chosen for the subgrid terms, making it unclear whether such models accurately capture the true physics.
@@ II.3 Diagnostics
Figure 3: Conservation of the mass and angular momentum integrals $M_{\textup{int}}$ and $J_{\textup{int}}$ for the representative SP2.54 case vs, retarded time $t_{\textup{ret}}:=t-r_{*}$, where $r_{*}$ is the extraction radius in tortoise coordinates $r_{*}=r+2M\ln(r/(2M)-1)$. The magenta lines denote the change in the surface integrals Eqs. (37) and (39) in Etienne et al. (2012a) evaluated on spherical surfaces at $r_{\textup{ext}}\simeq 273M$. The cyan line denotes the total EM energy (upper plot) and $z$ component of the total EM angular momentum (lower plot) that has passed across that surface. The green and red lines do the same for the fluid and GWs respectively. The black dashed line gives the total energy in the top panel, and total angular momentum in the bottom panel that should be conserved, adding in the lost energy and angular momentum.
@@ II.4 Criteria for a sGRB-compatible jet
As discussed in Sec.I, to be consistent with sGRB observations, the central engine needs to produce a an ultrarelativistic collimated jet with a maximum Lorentz factor of $\Gamma\gtrsim 20$ and, to match the kilonova observations, an ejecta mass of $\gtrsim 1\%$ of the total NSNS mass Coulter et al. (2017); Shappee et al. (2017); Arcavi et al. (2017). One highly non-trivial challenge in analyzing the result of GRMHD merger simulations is estimating the terminal Lorentz factor $\Gamma_{\infty}$, as our simulations have a finite spatial and temporal extent while we expect the velocity to increase with distance from the central engine (for instance, for a simple model of an ideal force-free paraboloidal jet the maximum Lorentz factor goes with distance along the jet $z$ as $\Gamma_{\textup{max}}\sim 0.3(z/M)^{1/2}$ Tchekhovskoy et al. (2008); McKinney and Blandford (2009); Narayan et al. (2010)). The physics of such jets, whether in the context of sGRB progenitors or active galactic nuclei, has been explored extensively through both numerical MHD and GRMHD studies (see e.g. Tchekhovskoy et al. (2008); McKinney and Blandford (2009); Narayan et al. (2010); Fernández et al. (2019); Nathanail et al. (2021); Gottlieb et al. (2022a); Mizuno (2022); Perucho and López-Miralles (2023) and the references cited therein) and semi-analytic studies (see e.g. Vlahakis and Konigl (2003); Beskin and Nokhrina (2005); Komissarov (2011); Lyubarsky (2011)) which make use of the assumptions of ideal MHD, special relativity, axisymmetry, steady state and radial self-similarity. For a steady state, axisymmetric ideal MHD flow integrating the equations of motion results in a number of conserved quantities along the poloidal component $\bm{B}_{p}$ of the magnetic field lines, which are parallel to the poloidal component of the fluid velocity $\bm{v}_{p}$ Vlahakis and Konigl (2003); Narayan et al. (2010); Lyubarsky (2011). One of these quantities is the ratio of total energy flux to rest-mass flux along a bundle of field lines, given by
$\mu:=$ energy fluxrest-mass $\frac{\textup{energy flux}}{\textup{rest-mass flux}}=\;\frac{\Gamma^{2}\rho_{0}hv_{p}+\tfrac{1}{4\pi}(\bm{E}\times\bm{B})\cdot\hat{\bm{v}}_{p}}{\Gamma\rho_{0}v_{p}},$
@@
Figure 4: Top panel: Pseudocolor plot of the quality factor $Q\equiv\lambda_{\textup{MRI}}/dx$ on the equatorial plane. Bottom panel: Rest-mass density normalised to its initial maximum value on a log scale along with $\lambda_{\textup{MRI}}$ (white line) on the meridional plane. The height of the line above the $x$ axis indicates the value of $\lambda_{\textup{MRI}}$ at $x=0$. Note that almost all of the area in the top image is red, indicating that the quality factor is $Q\gtrsim 10$ as required. We see that the height of the remnant and the height of the surrounding torus is larger than $\lambda_{\textup{MRI}}$ wavelength in all but a few regions, suggesting the MRI is active. The data shown here is for the SP2.57 case depicted at $\sim 17.8$ms post merger. However, it is representative for all cases, including the HMNS remnant in case IR2.70 prior to its collapse and the post-collapse accretion disk
@@ III Results
Figure 6: 2D plots in the meridional plane at 20ms post-merger for five of the simulations in Table 2. We show the rest-mass density $\rho_{0}$ on a log scale relative to the initial maximum rest-mass density (top row), $b^{2}/(2\rho_{0})$ which approximately corresponds to the magnetization $\sigma$ (second row), the Lorentz factor (third row), and the strength of the poloidal and toroidal magnetic field (fourth and fifth rows), with the magnetic field lines shown in the poloidal field plots. The scale on the two axes is the same. The arrows in the Lorentz factor plot (third row) indicate the flow velocities, and the arrow labelled on the left-hand side indicates the magnitude of the speed of light.
@@
Figure 7: Angular distribution of electromagnetic Poynting luminosity per solid angle for three cases at a distance $r=1024$km and time $t-t_{\textup{merge}}\sim 17$ms for $\theta\leq 30^{\circ}$, where $\theta$ is the polar angle from the $z$ axis (here “st” denotes “steradian”).
@@
Accretion into the BH down the polar axis and the acceleration of gas outwards due to the magnetic forces produces a low density ($\rho_{0}\;\lesssim 10^{7}\textup{g}\,\textup{cm}^{-3}$) evacuated funnel [^5] Notice that in our previous NSNS merger simulations where neutrinos were incorporated using the truncated-moment (M1) formalism the baryon pollution in the funnel was a factor of $\sim 10$ lighter than in this case Sun et al. (2022). with a half-opening angle at the base, defined in terms of the angular width of the magnetically dominated region, of $\sim 30^{\circ}$ (see the top left image of Fig. 6). The half-opening angle inferred from distant electromagnetic observations is significantly smaller, as at large distances the outflow approaches cylindrical flow Vlahakis and Konigl (2003), with an angle $\sim 10^{\circ}$ at distance $r=1024$km (see Fig. 7 left panel).
Figure 8: The accretion rate onto the BH remnant for the IR2.70 case (top), the fraction of the total rest mass outside the apparent horizon (middle), and the jet launching efficiency ($\eta=L/\dot{M}$ where $L=L_{\textup{EM}}+L_{\textup{fluid}}$ is the total jet luminosity) on log scales. The accretion rate decays as $\dot{M}\propto t^{-2}$ in the first $\sim 0.5$ms post-BH formation, as expected for highly magnetized disks Gottlieb et al. (2023), then rapidly transitions to $\dot{M}\propto t^{-1}$ as for hydrodynamically-dominated accretion, where accretion is driven by shocks and spirals in the disk Gottlieb et al. (2023). We also show the expected accretion efficiency for a Novikov-Thorne thin disk model (green dashed line, lower panel).
The low density and strong $\sim 10^{16}$ magnetic field enables a force-free environment with $b^{2}/(2\rho_{0})\sim\sigma\gg 1$ in the funnel above the BH poles, and the helical field enables the acceleration of an unbound, collimated, mildly relativistic ($\Gamma\sim 2.0$ in our simulation box) outflow, mainly along the outer regions of the funnel. This behaviour is consistent with the standard picture of Poynting-flux driven outflow reported in the literature (e.g. McKinney and Blandford (2009)) which we describe as an incipient jet. The accretion rate (see Fig. 8 top panel) shows a power-law decay with time, initially falling with $\dot{M}\propto t^{-2}$ in the first $\sim 0.5$ms, consistent with the expectation for a magnetically dominated disk Gottlieb et al. (2023), then rapidly transitions to $\dot{M}\propto t^{-1}$ for the remaining simulation time, consistent with a hydrodynamically-dominated disk where accretion is driven by shocks and density spirals Gottlieb et al. (2023). The accretion efficiency (see Fig. 8 bottom panel), $\eta:=(L_{\textup{fluid}}+L_{\textup{EM}})/\dot{M}$, increases to $\sim 0.09$, close to the value expected for the thin disk Novikov-Thorne model (see Shapiro (2005)), which has shown remarkably close agreement with numerical MHD simulations of BZ jets McKinney and Gammie (2004); Shapiro (2005), but below the maximum value for a magnetically-arrested disk (MAD) Lowell et al. (2024). From the $\dot{M}\propto t^{-1}$ accretion rate we can estimate the lifetime of the accretion disk as $\sim 2.2$s, very close to the typical lifetime of a sGRB, although our lack of knowledge of the precise details of how the energy from the incipient jet is converted to the $\gamma$-ray signal we observe limits the usefulness of this comparison.
@@
where $B_{\textup{pol}}$ is the poloidal magnetic field measured at the BH pole. The luminosity from Eq. (23) is plotted as connected dots in the middle panel of Fig. 9. We can see that the EM luminosity for the IR2.70 case converges towards the $L_{\textup{BZ}}$ value towards the end of our simulation as the additional power from the magnetized torus dies down and the BZ mechanism dominates. The EM luminosity magnitude lies within the narrow “universal range” of $10^{52\pm 1}$erg s-1 derived in Shapiro (2017) for accreting BH + disk systems resulting from compact binary mergers containing NSs, as well as from the magnetorotational collapse of massive stars. This, combined with the other criteria already validated for the very similar case reported in Ruiz et al. 2021 Ruiz et al. (2021), demonstrates that the BZ mechanism is operating in our simulation, as we concluded in Ruiz et al. (2016, 2018a); Paschalidis et al. (2015); Sun et al. (2022). One can relate the total electromagnetic luminosity $L_{\textup{EM}}$ observed in our simulations to the observed isotropic-equivalent $\gamma-$ray luminosity for a jet observed head-on as
$L_{\gamma,\textup{iso}}=\frac{1}{C_{\textup{col}}}\eta_{\textup{EM}}L_{\textup{EM}},$ (24)
where $C_{\textup{col}}$ is a factor that accounts for $\gamma$-ray collimation Khan et al. (2018), and $\eta_{\textup{EM}}$ corresponds to the efficiency of converting the outgoing Poynting flux in the simulation region to $\gamma-$ray photons in the emission zone. For perfect collimation, where all the emission is contained within a homogenous jet of half-opening angle $\theta$, we have $C_{\textup{col}}=1-\cos\theta\sim 10^{-2}-10^{-1}$ for $\theta\sim 10^{\circ}-30^{\circ}$ Escorial et al. (2023).
The prompt emission mechanism is still an open question Parsotan and Ito (2022) making the efficiencies highly uncertain Beniamini et al. (2016), although $\sim 10^{-1}$ has been taken as a fiducial value Khan et al. (2018) for the EM Poynting luminosity for outflow from a black hole central engine. For our cases with a SMNS central engine we see indications of mixing between the central funnel and the debris torus, associated energy loss from the outflow, and larger baryon pollution compared to the black hole central engine case, which may indicate that the efficiency in these cases is $\ll 10^{-1}$. Taking $\eta_{\textup{EM}}\sim 10^{-2}-10^{-1}$ we get an estimated $L_{\gamma,\textup{iso}}\sim 10^{52}-10^{54}$ erg s-1. The lower end of this estimate is compatible with sGRB observations Li et al. (2016); Shapiro (2017); Beniamini et al. (2020). With regards to the higher end of the $L_{\gamma,\textup{iso}}$ estimate, it is possible that the large luminosity is due to the artificially strong magnetic field used, or the choice of magnetic field geometry, although as discussed in Sec. I, previous works have found that a strong large-scale poloidal magnetic field is a requirement for jet launching Ruiz et al. (2020a); Etienne et al. (2012b). Hence, our simulations represent a “best-case” scenario for jet formation. We intend to investigate the impact of magnetic field strength and topology in a future work.
Although we only observe a mildly relativistic outflow within our simulation region, an extensive literature suggests that the BZ effect alone can ultimately accelerate gas to ultrarelativistic ($\Gamma_{\infty}\gg 100$) velocities, even with significant baryon pollution (see e.g. Paschalidis et al. (2015); McKinney (2005)) and therefore provide the central engine for a sGRB-compatible jet.
Figure 9: Luminosity via fluid outflow, EM radiation and gravitational waves, vs. retarded time since $t_{\textup{merge}}$ (defined as the time of peak GW emission) extracted over a sphere at $r_{\textup{ext}}=1024\textup{km}\sim 260M$. The purple dots and connecting line in the middle plot show the expected BZ electromagnetic luminosity from Eq. (23) for the BH mass, spin and approximate $B_{\textup{pol}}$ for case IR2.70.